# Long-time soliton dynamics via a coarse-grained space-time method

### v2

@@ Appendix C Boundary conditions in DEC
Ordinary Neumann boundary condition: examples of this type of boundary condition were presented in Eqs. 17, 18, and 49, in which the conditions on $\partial_{x}\varphi$ at the boundary incorporate how external magnetic field and injected boundary currents affect the internal dynamics of the junction. These BCs, which set the value for $\phi_{x}$ at the boundary, can be directly implemented on the boundary edges $e_{x}[(1,j),(2,j)]$ and $e_{x}[(N_{x}-1,j),(N_{x},j)]$ at every time step $j$ in the spacetime grid (Fig. 20). Similarly, boundary conditions imposed solely on the time derivative of the scalar field, $\partial_{t}\varphi$, can be implemented on the field $\phi_{t}$ on the boundary edges $e_{t}[(1,j),(1,j+1)]$ and $e_{t}[(N_{x},j),(N_{x},j+1)]$.
@@
[…]
@@
$ik=\pm i\omega\sqrt{1-\frac{U(x,t)}{\omega^{2}}}\approx\pm i\omega\left[1+\frac{U(x,t)}{2(i\omega)^{2}}+...\right]$ (43)
@@
Similar outgoing conditions for the right boundary are obtained by flipping the sign on the left-hand side of Eqs. 44-45. Higher order boundary conditions can be readily obtained by keeping more terms in the expansion in Eq. 43.
@@
$+\left(\Delta t/\Delta x\right)\left\{\phi_{x}[(1,j),(2,j)]-\phi[(1,j\!-\!1),(2,j\!-\!1)]\right\}$
@@ Appendix D Profiling simulation times
A comparison on the runtimes of DEC-QED and the Euler method is presented in Fig. 21. The problem used for the comparison is the dynamics of a sine-Gordon fluxon trapped inside an unperturbed Josephson junction of length $L=100$. The simulated fluxon moves inside the junction at speed $u=0.55$ and for a duration of $T_{m}=50000$. For all the spatial step size $\Delta x$ tested, we choose $\Delta t=0.8\Delta x$. Since the two methods have the same complexity, the number of computations scales linearly with the number of gridpoints in the spacetime grid. However, because DEC-QED deals with the first-order PDE of the edge fields, it is faster than solving the second-order noninear wave equation for the scalar field $\varphi$ using the traditional Euler method. As a result, DEC-QED is slightly faster than the Euler method regardless of the step size used in the simulation.
@@ Appendix E Modification of vortex dynamics due to biased boundary conditions
Figs. 22a, 22b, and 22c show the results for $\varphi$, $\phi_{t}$, and $\phi_{x}$ with $\eta=0.002$ and $\xi=0.006$. We observe a gradual speeding up of the soliton over time. This is because every time the soliton collides with a boundary where $\eta$ and $\xi$ are non-zero, it is in contact with an externally injected current and a magnetic field that either add to or decrease its kinetic energy depending on the relative orientation of the current of the (anti)vortex with these boundary effects. For the parameters chosen for the simulations in Figs. 22a-c, after each round-trip (i.e. after colliding with the boundary on both ends) the soliton obtains a small additional energy and speeds up. However, the acceleration decreases over time as the circulating current of the (anti)vortex approaches the same velocity as the boundary current. In total, the dynamics shown in Fig. 22a-c has 453 collisions with the boundaries, covering a total distance slightly greater than 18875 times its original size. To compare with the vortex dynamics in Figs. 4 in the main text, where we have the same initial conditions and parameters except that $\eta=\xi=0$, the fluxon only collides 278 times with the boundary.
@@ Appendix F Vortex in narrow constrictions
Typically, in a long Josephson junction, it is difficult to maintain uniform thickness and material properties throughout the entire length of the junction. The microshorts discussed in the previous section are extreme cases, where the thickness of the insulating region is near zero and the superconducting islands are essentially in contact with each other. In general, the thickness of the insulating region can vary along the junction, and the sine-Gordon equation can be modified accordingly to capture this behavior
@@
where the critical current $\mu(x)$ is now a function of $x$. Using DEC-QED, we consider the evolution of a fluxon in a junction that has a narrow constriction inside (see Fig. 23a), where the insulating region is thinner than the rest of the junction, leading to a stronger critical current there. In general, the critical current depends exponentially on the thickness of the insulating region. In the limit of thin junctions, the dependence is approximately inversely proportional to the thickness of the junction, $a$,  [43], i.e. $\mu\propto 1/\sinh(\xi a)\sim 1/a$. In Fig. 23b-c the dynamics of a fluxon moving in a constricted junction is shown. The junction length is $L=200$, the length of the constriction is $\ell=40$, and the lengths of the two tapered regions that connect the constriction with the rest of the junction are $b=10$. In Fig. 23b, the critical current inside the constriction is $\mu_{cs}=10$, while the critical current at the normal regions are unity. The fluxon has an initial velocity of $u=0.85$ and travels without resistive loss. Similar to the microshort case, this narrow constriction draws energy from the fluxon, preventing it from entering and repeling it from the constriction. The movement of this fluxon is therefore limited to bouncing between the junction boundary and the edge of the constriction.
If the critical current in the constriction is sufficiently relaxed, which is the situation in Fig. 23c where $\mu_{cs}=3$, the fluxon can enter the narrow channel. When the fluxon is inside this potential barrier, part of the kinetic energy is transferred to potential energy and the soliton slows down, as can be seen by the change in the slope of the trace in Fig. 23c. Note that the higher critical current in a narrow junction means the Swihart velocity is also lower there. When the soliton has just entered the constriction, however, it can still have a velocity higher than the maximum velocity allowed in this region. As the soliton is abruptly slowed down to below the Swihart velocity in the constriction, it also releases energy in the form of radiation. These are similar to Cherenkov radiations that appear in nonlinear optical systems that have high-order dispersions such as in fiber optics [67]. This effect is evidently seen in Fig. 23e, where the dynamics of a soliton moving through a series of three constrictions is shown. The fringes caused by radiation coming out of the soliton are most intense at the bending corners of the trajectory, where the soliton enters a narrow constriction and has to slow down. In this example, after the soliton exits the last constriction, roughly $\approx 3\%$ of its initial kinetic energy has been transferred to radiations.
@@ Appendix G Creation of solitons using applied external pulse
So far we have discussed a wide variety of possible situations involving solitons embedded in Josephson junctions and demonstrated how DEC-QED is able to reliably capture the evolution of these systems. These simulations all begin with solitons as initial conditions. However, in a realistic experimental setup, these fluxons or breathers have to be created by a well-controlled application of external fields [79, 80]. Therefore, it is also important to be able to simulate how these solitons can emerge from the background via external excitation. In this section, we discuss how DEC-QED is a suitable tool for this.
We consider a long junction subjected to an external magnetic pulse at one boundary edge of the junction. This pulse affects the junction through a boundary condition of the form
@@
When a junction experiences a finite pulse at the boundary, radiations, fluxons, and antifluxons of any velocity $u\leq 1$ can occur on that boundary and travel into the junction. This is demonstrated in Fig. 24b, where we choose a junction whose length is $L=160$. The pulse has a total duration of $T_{m}=250$, with frequency $\omega=0.8$. The rise time and fall time are $\sigma_{rise}=\sigma_{fall}=10$, and the pulse amplitude after the rise time and before termination is $A=1.5$. In the 2D $x-t$ heatmap in Fig. 24b one can see very clearly that the generated excitations are a series of fluxons, antifluxons, and breathers, all moving at different speeds and interfering with each other, plus some background radiations. It is also worthwhile to note that instead of looking at this holistic view of the full dynamics, one can plot the field at a specific time slice (as done in Fig. 24 where $\phi_{x}$ at $t=160$ is shown), and it appears as if there is no structure to the dynamics generated by the pulse which is actually composed of mostly solitons.
Given the efficiency and accuracy of the numerical scheme, one can easily sweep a large parameter space to tune the pulse to produce the desired results. To demonstrate, in Fig. 24c and 24d we plot the $\phi_{x}$ and $\varphi$ fields for a setup in which the external pulse creates exactly one fluxon, one anti-fluxion, and one breather. To find the correct pulse, we had sampled the parameter space of $(A,\omega,\sigma_{rise},\sigma_{fall},T_{p})$ in the following ranges and steps: $21\!\leq\!A\!\leq\!$ 2 with step $\Delta A\!=\!0.1$, $155\!\leq\!\sigma_{rise}\!=\!\sigma_{fall}\!\leq\!$ 15 with step $\Delta\sigma\!=\!1$, $0.5\leq\omega\leq 0.9$ with step $\Delta\omega\!=\!0.1$, and $8040\!\leq\!T_{p}\!\leq\!$ 80 with $\Delta T_{p}\!=\!2$. This sampling amounts to 12700 coordinates in the parameter space. We choose a junction length of $L=160$ that is discretized by grid spacings $\Delta x=0.02$, while the simulated time is $T_{m}=250$ with $\Delta t=0.016$. For each set of parameters for the pulse, the simulation is therefore done with a spacetime mesh of $125$ million grid points. The total time it takes for the sweep to return the desired pulse is $5.5$ hours.
@@ Appendix H Massive capacitor discharge
Figure 25: Time evolution of the electric field due to a sharged capacitor and the back reaction of the fermionic fluid. The capacitor is charged at $Q=100$, with the plates located at ±$\pm L/2=\pm 20$, and $g=1.2$. The dynamics is captured up to $T_{m}=2000$. The electric field $E$ at $x\!=\!0$ is shown in (a), (b), and (c) for when the mass of the fermions are $m=0,\ m=0.1,$ and $m=1.0$, respectively. In a similar order, the evolution of $E$ at $x\!=\!100$ when $m=0,\ m=0.1,$ and $m=1.0$ are shown in (d), (e) and (f).
[…]
@@ Appendix I Different long-time scenarios of the Schwinger soliton-antisoliton pair
Figure 26: Possible dynamics of a system with large effective mass ($g=0.4$) and is initially composed of a soliton-antisoliton pair moving towards each other at the speed $u=0.55$. The evolution over time of the energies stored inside a finite domain $-16\leq x\leq 16$ are shown in (a), (b), (c) and (d) for $d=22,\ d=16.5,\ d=13.75,$and $d=11$, in that order.
@@
When the long-time configuration around the center is a pair of secondary breathers oscillating around each other (resulting in a steady beating pattern with the meson density periodically dispersing and refocusing), the energy evolution also exhibits the beating behavior (see Fig. 26a). The close-up inset reveals that, in the steady state, the energy associated with the quadratic potential term, $(\partial_{x}\varphi)^{2}$ (blue curve), undergoes periodic collapse and revival as its dynamics intermittently synchronize and desynchronize with the other energy components. At steady state, the background meson oscillations near the system’s center of mass also settle into a standing wave pattern. This indicates the absence of propagation in these oscillations, meaning that no energy escapes from the finite region surrounding the core. The energy dynamics in Fig. 26b corresponds to scenario shown in Fig. 17b in which the late-time dynamics is a single breather remaining at $x\!=\!0$; the total energy in the region around the origin reaches a steady state, with energy cycling continuously at a fixed frequency among kinetic, potential, and electric field contributions. Lastly, in Figs. 26c-d we show the evolution of the energy terms over time when the initial solitons turn into breathers and scatter off to infinity without generating further mesons in the region between them. The total energy within the finite region around $x=0$ drops to zero and remains there once the breathers have left.

### v3

@@ Appendix C Boundary conditions in DEC
Ordinary Neumann boundary condition: examples of this type of boundary condition were presented in Eqs. 18, 19, and 52, in which the conditions on $\partial_{x}\varphi$ at the boundary incorporate how external magnetic field and injected boundary currents affect the internal dynamics of the junction. These BCs, which set the value for $\phi_{x}$ at the boundary, can be directly implemented on the boundary edges $e_{x}[(1,j),(2,j)]$ and $e_{x}[(N_{x}-1,j),(N_{x},j)]$ at every time step $j$ in the spacetime grid (Fig. 20). Similarly, boundary conditions imposed solely on the time derivative of the scalar field, $\partial_{t}\varphi$, can be implemented on the field $\phi_{t}$ on the boundary edges $e_{t}[(1,j),(1,j+1)]$ and $e_{t}[(N_{x},j),(N_{x},j+1)]$.
@@
with $U(x,t)\ll 1$. By performing a Fourier transform $\varphi(x,t)\!=\!1/\!(4\pi^{2})\!\int\!\!\int\!e^{i(kx+\omega t)}\hat{\varphi}(k,\omega)dkd\omega$ to Eq. 44 and collecting the terms in the integrand we obtain
$k^{2}$ $=\omega^{2}-\frac{1}{4\pi^{2}}\!\int\!\int\!\hat{U}(k^{\prime},\omega^{\prime})e^{i(k^{\prime}x+\omega^{\prime}t)}dk^{\prime}d\omega^{\prime}$
@@
[…]
@@
Similar outgoing conditions for the right boundary are obtained by flipping the sign on the left-hand side of Eqs. 47-48. Higher order boundary conditions can be readily obtained by keeping more terms in the expansion in Eq. 46.
@@
$+\left(\Delta t/\Delta x\right)\left\{\phi_{x}[(1,j),(2,j)]-\phi_{x}[(1,j\!-\!1),(2,j\!-\!1)]\right\}$
@@ Appendix D Profiling simulation times
A comparison on the runtimes of DEC-QED and the Euler method is presented in Fig. 21. The problem used for the comparison is the dynamics of a sine-Gordon fluxon trapped inside an unperturbed Josephson junction of length $L=100$. The simulated fluxon moves inside the junction at speed $u=0.55$ and for a duration of $T_{m}=50000$. For all the spatial step size $\Delta x$ tested, we choose $\Delta t=0.8\Delta x$. Since the two methods have the same complexity, the number of computations scales linearly with the number of gridpoints in the spacetime grid. However, because DEC-QED deals with the first-order PDE of the edge fields, it is faster than solving the second-order noninear wave equation for the scalar field $\varphi$ using the traditional Euler method. As a result, DEC-QED is slightly faster than the Euler method regardless of the step size used in the simulation. One should also note that if the sine term is removed and the dynamcis become linear, the two methods will have the same convergence rate for the derived fields $\phi_{x}$ and $\phi_{t}$. By nature of it being a local and time-propagating method DEC-QED also requires that the Courant condition is respected in order for the evolution to be stable. In a nonlinear and out-of-equilibrium setting, however, DEC-QED remains stable for a considerably longer simulations compared to standard Euler’s method - as shown in Fig. 4 and Fig. 15.
@@ Appendix E Modification of vortex dynamics due to biased boundary conditions
Figs. 22a, 22b, and 22c show the results for $\varphi$, $\phi_{t}$, and $\phi_{x}$ with $\eta=0.002$ and $\xi=0.006$. We observe a gradual speeding up of the soliton over time. This is because every time the soliton collides with a boundary where $\eta$ and $\xi$ are non-zero, it is in contact with an externally injected current and a magnetic field that either adds to or decreases its kinetic energy depending on the relative orientation of the current of the (anti)vortex with these boundary effects. For the parameters chosen for the simulations in Figs. 22a-c, after each round-trip (i.e. after colliding with the boundary on both ends) the soliton obtains a small additional energy and speeds up. With each subsequent round trip, the velocity of the (anti)vortex approaches that of the boundary current. As this happens, the (anti)vortex speeds up by less and less with each collision. In total, the dynamics shown in Fig. 22a-c has 453 collisions with the boundaries, covering a total distance slightly greater than 18875 times its original size. To compare with the vortex dynamics in Figs. 4 in the main text, where we have the same initial conditions and parameters except that $\eta=\xi=0$, the fluxon only collides 278 times with the boundary.
@@ Appendix F Vortex in narrow constrictions
Typically, in a long Josephson junction, it is difficult to maintain uniform thickness and material properties throughout the entire length of the junction. The microshorts discussed in Section V.2.3 are extreme cases, where the thickness of the insulating region is near zero and the superconducting islands are essentially in contact with each other. In general, the thickness of the insulating region can vary along the junction, and the sine-Gordon equation can be modified accordingly to capture this behavior
@@
where the critical current $\mu(x)$ is now a function of $x$. Using DEC-QED, we consider the evolution of a fluxon in a junction that has a narrow constriction inside (see Fig. 23a), where the insulating region is thinner than the rest of the junction, leading to a stronger critical current there. In general, the critical current depends exponentially on the thickness of the insulating region. In the limit of thin junctions, the dependence is approximately inversely proportional to the thickness of the junction, $a$, [dec-qed], i.e. $\mu\propto 1/\sinh(\xi a)\sim 1/a$. In Fig. 23b-c the dynamics of a fluxon moving in a constricted junction is shown. The junction length is $L=200$, the length of the constriction is $\ell=40$, and the lengths of the two tapered regions that connect the constriction with the rest of the junction are $b=10$. In Fig. 23b, the critical current inside the constriction is $\mu_{cs}=10$, while the critical current at the normal regions are unity. The fluxon has an initial velocity of $u=0.85$ and travels without resistive loss. Similar to the microshort case, this narrow constriction draws energy from the fluxon, preventing it from entering and repeling it from the constriction. The movement of this fluxon is therefore limited to bouncing between the junction boundary and the edge of the constriction.
If the critical current in the constriction is sufficiently relaxed, which is the situation in Fig. 23c where $\mu_{cs}=3$, the fluxon can enter the narrow channel. When the fluxon is inside this potential barrier, part of the kinetic energy is transferred to potential energy and the soliton slows down, as can be seen by the change in the slope of the trace in Fig. 23c. Note that the higher critical current in a narrow junction means the Swihart velocity is also lower there. When the soliton has just entered the constriction, however, it can still have a velocity higher than the maximum velocity allowed in this region. As the soliton is abruptly slowed down to below the Swihart velocity in the constriction, it also releases energy in the form of radiation. These are similar to Cherenkov radiations that appear in nonlinear optical systems that have high-order dispersions such as in fiber optics [akhmediev1995cherenkov]. This effect is evidently seen in Fig. 23e, where the dynamics of a soliton moving through a series of three constrictions is shown. The fringes caused by radiation coming out of the soliton are most intense at the bending corners of the trajectory, where the soliton enters a narrow constriction and has to slow down. In this example, after the soliton exits the last constriction, roughly $\approx 3\%$ of its initial kinetic energy has been transferred to radiations.
@@ Appendix G Creation of solitons using applied external pulse
So far we have discussed a wide variety of possible situations involving solitons embedded in Josephson junctions and demonstrated how DEC-QED is able to reliably capture the evolution of these systems. These simulations all begin with solitons as initial conditions. However, in a realistic experimental setup, these fluxons or breathers have to be created by a well-controlled application of external fields [kivshar1991solitoncreation, sakai1987solitoncreation]. Therefore, it is also important to be able to simulate how these solitons can emerge from the background via external excitation. In this section, we discuss how DEC-QED is a suitable tool for this.
We consider a long junction subjected to an external magnetic pulse at one boundary edge of the junction. The mechanism of generating solitons by applying a boundary pulse is similar to those studied in the physics of nonlinear supratransmission [geniet2002energy, khomeriki2009fluxon, schlawin2017terahertz, desantis2022generation, desantis2022supratransmission]. In our current setting, the effect of the pulse to the junction is modeled through a boundary condition of the form
@@
When a junction experiences a finite pulse at the boundary, radiations, fluxons, and antifluxons of any velocity $u\leq 1$ can occur on that boundary and travel into the junction. This is demonstrated in Fig. 24b, where we choose a junction whose length is $L=160$. The pulse has a total duration of $T_{m}=250$, with frequency $\omega=0.8$. The rise time and fall time are $\sigma_{rise}=\sigma_{fall}=10$, and the pulse amplitude after the rise time and before termination is $A=1.5$. In the 2D xtx-t heatmap in Fig. 24b one can see very clearly that the generated excitations are a series of fluxons, antifluxons, and breathers, all moving at different speeds and interfering with each other, plus some background radiations. It is also worthwhile to note that instead of looking at this holistic view of the full dynamics, one can plot the field at a specific time slice (as done in Fig. 24 where $\phi_{x}$ at $t=160$ is shown), and it appears as if there is no structure to the dynamics generated by the pulse which is actually composed of mostly solitons.
Given the efficiency and accuracy of the numerical scheme, one can easily sweep a large parameter space to tune the pulse to produce the desired results. To demonstrate, in Fig. 24c and 24d we plot the $\phi_{x}$ and $\varphi$ fields for a setup in which the external pulse creates exactly one fluxon, one anti-fluxion, and one breather. To find the correct pulse, we had sampled the parameter space of $(A,\omega,\sigma_{rise},\sigma_{fall},T_{p})$ in the following ranges and steps: $1\!\leq\!A\!\leq\!2$ with step $\Delta A\!=\!0.1$, $5\!\leq\!\sigma_{rise}\!=\!\sigma_{fall}\!\leq\!15$ with step $\Delta\sigma\!=\!1$, $0.5\leq\omega\leq 0.9$ with step $\Delta\omega\!=\!0.1$, and $40\!\leq\!T_{p}\!\leq\!80$ with $\Delta T_{p}\!=\!2$. This sampling amounts to 12700 coordinates in the parameter space. We choose a junction length of $L=160$ that is discretized by grid spacings $\Delta x=0.02$, while the simulated time is $T_{m}=250$ with $\Delta t=0.016$. For each set of parameters for the pulse, the simulation is therefore done with a spacetime mesh of $125$ million grid points. The total time it takes for the sweep to return the desired pulse is $5.5$ hours.
@@ Appendix H Massive capacitor discharge
Figure 25: Time evolution of the electric field due to a sharged capacitor and the back reaction of the fermionic fluid. The capacitor is charged at $Q=100$, with the plates located at $\pm L/2=\pm 20$, and $g=1.2$. The dynamics is captured up to $T_{m}=2000$. The electric field $E$ at x=0x\!=\!0 is shown in (a), (b), and (c) for when the mass of the fermions are m=0,m=0.1,m=0,\ m=0.1,\ and $m=1.0$, respectively. In a similar order, the evolution of $E$ at x=100x\!=\!100 when m=0,m=0.1,m=0,\ m=0.1,\ and $m=1.0$ are shown in (d), (e) and (f).
In Section VI.1 we studied the backreaction dynamics of massless fermions under the influence of a charged capacitor. In this appendix, we present the results for the capacitor discharge dynamics when the fermions are massive. Figs. 25a-c provide direct comparisons of the electric field at the center of the capacitor when the fermion mass is m=0, 0.1,m=0,\ 0.1, and $1.0$, respectively. Figs. 25d-f present the electric field at a finite distance away from the capacitor. As evidently shown in Figs. 25, a finite mass $m$ leads to fast oscillations in the field due to the presence of the term $\sin\varphi$ in the equation of motion, with larger $m$ resulting in oscillations with higher amplitudes.
@@ Appendix I Different long-time scenarios of the Schwinger soliton-antisoliton pair
Figure 26: Possible dynamics of a system with large effective mass ($g=0.4$) and is initially composed of a soliton-antisoliton pair moving towards each other at the speed $u=0.55$. The evolution over time of the energies stored inside a finite domain $16\leq x\leq 16$ are shown in (a), (b), (c) and (d) for d=22,d=16.5,d=13.75,d=22,\ d=16.5,\ d=13.75,\ and $d=11$, in that order.
@@
When the long-time configuration around the center is a pair of secondary breathers oscillating around each other (resulting in a steady beating pattern with the meson density periodically dispersing and refocusing), the energy evolution also exhibits the beating behavior (see Fig. 26a). The close-up inset reveals that, in the steady state, the energy associated with the quadratic potential term, $(\partial_{x}\varphi)^{2}$ (blue curve), undergoes periodic collapse and revival as its dynamics intermittently synchronize and desynchronize with the other energy components. At steady state, the background meson oscillations near the system’s center of mass also settle into a standing wave pattern. This indicates the absence of propagation in these oscillations, meaning that no energy escapes from the finite region surrounding the core. The energy dynamics in Fig. 26b corresponds to scenario shown in Fig. 17b in which the late-time dynamics is a single breather remaining at x=0x\!=\!0; the total energy in the region around the origin reaches a steady state, with energy cycling continuously at a fixed frequency among kinetic, potential, and electric field contributions. Lastly, in Figs. 26c-d we show the evolution of the energy terms over time when the initial solitons turn into breathers and scatter off to infinity without generating further mesons in the region between them. The total energy within the finite region around $x=0$ drops to zero and remains there once the breathers have left.